OpenStudy (anonymous):

What is the perimeter of a parallelogram with sides of 5 and 7?

OpenStudy (anonymous):

The other two sides will also be 5 and 7, So the total of two 5's and two 7's is 24 units.

OpenStudy (anonymous):

p = 2a+2b p = 2(5)+2(7) p = 10+14 p = 24

jhonyy9 (jhonyy9):

p=24